OpenStudy (anonymous):

9^10 x 27^20 = ??? can you give me the explanation please ! this is my retrice

OpenStudy (anonymous):

I can give an answer but not so well of an explanation.

OpenStudy (anonymous):

why ?

OpenStudy (anonymous):

Because Math questions I get from Mathway.com, but if you want and explanation I would try looking at Khan academy they give amazing explanations.

OpenStudy (anonymous):

ok. thanks :) but what is your answer?

OpenStudy (anonymous):

Well, If it is 9^10 x 27^20 it can not be calculated. https://mathway.com/

OpenStudy (anonymous):

why ? :(

OpenStudy (anonymous):

can you explain with me why i cant simplify this :(

Miracrown (miracrown):

You could just type it into you calculator and get an answer, or we can mathematically break it down into smaller, more manageable parts than into a whole

Miracrown (miracrown):

When you insert this into the calc: \[9^{10} \space \times 27^{20} \space = 1.4780883e+38\]

Miracrown (miracrown):

the E+38 means go 38 decimal places to the right

Miracrown (miracrown):

We can break it down as well, I'll help you with the start ... We want to break 9 down into parts, we can do this as such 9 is 3 times 3 or 3^2

Miracrown (miracrown):

So how could we do this for 27? \[(3^2)^10*(?)^20 \]

Miracrown (miracrown):

Where the question mark is where 27 is so how can we then proceed to break that down into parts of maybe 3? @Loonie

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

what is the break parts of 27 ?

OpenStudy (anonymous):

its (3)^3

OpenStudy (mrnood):

whoaaa - do not rush to put numbers on this 9 = 3^2 27 = 3^3 Us the properties of indices to simplify the expression to a simple 3^x

OpenStudy (anonymous):

Nicely done MrNood

OpenStudy (anonymous):

how can i explain to my prof ?